How do you solve for Q in this equation?
0=-1400000+(60*Q/.16)

Answer by Alan3354(69443)

How do you solve for Q in this equation?
0=-1400000+(60*Q/.16)
------------------
Same as always, get it on one side of the equal sign by itself.
0=-1400000+(60*Q/.16)
Add 1400000 to both sides
1400000 = 60*Q/.16
Multiply both sides by .16
224000 = 60*Q
Divide by 60
Q = 224000/60
Q = 3733 1/3